OpenStudy (anonymous):

Need Help in midterm q1 (I1,I2 Power ,last one)

OpenStudy (anonymous):

my values The elements in this circuit have the following values: R1=100Ω, R2=100Ω, R3=200Ω, V1=5V, I1=0.01A and I2=0.001A

OpenStudy (anonymous):

powr of I1=vb*I1 n for I2 it is -Vb*I2...try...

OpenStudy (anonymous):

@rakhiraveendran ans not comming

OpenStudy (anonymous):

your Vb is correct??? i got it correct...!!!

OpenStudy (anonymous):

P_I1 = V1 * I1 or vb * i1.. and here what is vb

OpenStudy (anonymous):

vb is voltage at node b...the answer to the second question

OpenStudy (anonymous):

i have tried 0.05,0.042 for last 2nd and -0.005, -0.00042

OpenStudy (anonymous):

@hrene: if u got the anser for VB,den simply substitute in the equation...den u'll get power...or shud v solve from the beginning?

OpenStudy (anonymous):

I1 = 0.034 i2 = -0.0034

OpenStudy (anonymous):

One way is very simple. Just use Circuit Sandbox . There place the your circuit wit own values of components, and then you will get A and B points voltages . And then calculate power is easy. Power on R1 is (V1-A)^2/R1; Power on R2 is (A-B)^2/R2; If R1=R2 - power on these resistors is equal.. Power on R3 is B^2/R3; Power from V1: calculated current from V1*V1 Power from I1: B*I1 Power from I2: -B*I2